OpenStudy (anonymous):

Solve the equation. 9 - w = 19

OpenStudy (diyadiya):

subtract 9 from both sides

OpenStudy (diyadiya):

9-w -9 =19-9 -w=10 divide both sides by -1 you'll get the answer

OpenStudy (anonymous):

okay thank you

OpenStudy (diyadiya):

yw